# Educating LLMs like Human Students: Structure-aware Injection of Domain Knowledge

### v1

@@ Abstract
This paper presents a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists. It significantly minimizes the training corpus requirement to a mere 0.3% while achieving an impressive 50% of traditional knowledge injection performance. Our method is inspired by the educational processes for human students, particularly how structured domain knowledge from textbooks is absorbed and then applied to tackle real-world challenges through specific exercises. Based on this, we propose a novel two-stage knowledge injection strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT). In the SCPT phase, we organize the training data into an auto-generated taxonomy of domain knowledge, enabling LLMs to effectively memorize textual segments linked to specific expertise within the taxonomy’s architecture. Subsequently, in the SSFT phase, we explicitly prompt models to reveal the underlying knowledge structure in their outputs, leveraging this structured domain insight to address practical problems adeptly. Our ultimate method has undergone extensive evaluations across model architectures and scales, using closed-book question-answering tasks on LongBench and MMedBench datasets. Remarkably, our method matches 50% of the improvement displayed by the state-of-the-art MMedLM2 on MMedBench, but with only 0.3% quantity of the training corpus. This breakthrough showcases the potential to scale up our StructTuning for stronger domain-specific LLMs. Code will be made public soon.
@@ 1 Introduction
Large language models (LLMs) have recently seen extensive deployment across various applications [42, 1, 20, 6]. When adapting foundational models (e.g., Llama series [39, 40, 2]) for specialized AI assistants in distinct domains [35, 13], developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) [25] and domain knowledge injection [14]. While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses[55, 8]. Thus, another avenue of research focuses on injecting new knowledge into models via training techniques [11, 18, 30].
Continual pre-training [37, 19] has been preferred for integrating new, domain-specific knowledge into existing LLMs [9, 44, 35]. Nevertheless, it often involves costly auto-regressive training on billions of tokens aimlessly curated from the internet rather than from precise domain-specific textbooks [21]. For example, MMedLM [35] curates 25.5B tokens to derive a medical model, and DeepSeek-Coder [13] uses 2T tokens for coding adaptation. The common failure to learn effectively from limited textbook content has been attributed to insufficient data diversity [58], which however violates the observation during the human education process in Fig. 1: students gain knowledge by sequentially studying from textbooks, reviewing knowledge points and structures, and applying this knowledge through proper exercises. In this process, all the new data to learn is textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts [23, 52].
@@
Inspired by this, we propose to inject the domain knowledge from textbooks into LLMs, like educating a human student, through a novel two-stage training strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
@@
In the SCPT stage, we argue that high-quality textbook data alone can adequately infuse initial domain knowledge [12], where the organization of training corpora is crucial. Traditionally, text corpora are merely concatenated and divided into chunks of 2048 [35] or 4096 [13], while the inherent structure of the texts (e.g., catalogs of textbooks) is disregarded. Instead, we propose an automatic approach to maintain each chunk’s knowledge structure. We view each chunk as a knowledge point, employing advanced LLMs to extract domain knowledge taxonomy from the corpus efficiently, bypassing the need for manual annotation. Subsequently, LLMs are trained to predict textual content under the condition of linked knowledge points within the taxonomy, integrating individual training chunks with the entire knowledge framework. Ultimately, models are asked to memorize the knowledge structure to review the whole domain knowledge system.
@@
In the SSFT stage, the goal shifts from knowledge injection to enabling LLMs to recall and utilize their acquired knowledge to tackle real-world challenges. We explicitly elicit knowledge structures in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses. To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by advanced LLMs such as GPT4 [1] or LLaMA3 [2]. In the scenarios with existing QA pairs like MMedBench [35], we retrieve related knowledge structures and content, instructing LLaMA3 to provide explanations from questions to answers based on these structures. For datasets lacking specific QA samples like LongBench [4], we randomly select knowledge branches from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for exercises.
@@
Our ultimate approach, termed StructTuning, outperforms conventional methods in domain knowledge injection by emulating human learning processes through SCPT and SSFT phases. We extensively evaluate StructTuning’s effectiveness across different model architectures and sizes. For domain-adapted language models, we first examine their capability to recall injected knowledge through open-ended QA on the LongBench [4] dataset, then assess their application of this knowledge in addressing real-world issues via multiple-choice QA on the MMedBench dataset [35]. Results from both evaluations underscore the superiority of StructTuning. Remarkably, our strategy achieves a 50% improvement in knowledge injection compared to MMedLM2, the current benchmark in the medical domain, using merely 0.3% of the training data requirement. These findings underscore the potential to scale our method for enhancing domain-specific AI assistants significantly.
@@
  • We conducted extensive investigations on our StructTuning strategy on various data and model settings, and achieved competitive knowledge injection performance on MMedBench only using 0.3% of the training corpus.
@@ 2 Related Work
Domain Adaptation for Large Language Models. While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications [56, 43]. Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content [37, 50], with parameter-efficient tuning methods devised to curtail training costs [18, 29]. To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape [30, 35]. Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like law [9, 32], finance [48, 26], medicine [44, 35], and coding [36, 13]. Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
Conditional Language Modeling. The idea of continual pre-training language models on domain corpus in the condition of the knowledge structure is mainly inspired by CTRL [22]. Keskar et al. [22] demonstrates the effectiveness of steering text generation through control codes (one or two words) that signify the desired genre, style, or task. In the era of LLM, system prompt plays a similar role in controlling models’ responses to adapt to different needs and functionalities, such as role-playing, language style transfer, task setting, and behavior setting [7, 45, 3]. Our SCPT approach extends the control codes or system prompts to domain-specific knowledge structures, so as to guide the learning process and tailor the model’s output more closely to specialized fields.
Data Augmentation and Synthesis. Due to the lack of high-quality datasets, data augmentation has emerged as a promising solution to mimic real-world patterns [28]. Traditionally, these methods aim to artificially expand the training dataset size [49, 31] or generate entirely new samples that could help models learn better or adapt to specific tasks [38]. Yet, such methods often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity [33, 30], leading to potentially suboptimal training outcomes when applied for domain adaptations [30, 38]. Unlike these traditional approaches, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
@@ 3 Methodology
This section presents our methodology to inject domain knowledge into pre-trained LLMs by utilizing the inherent knowledge structure, and Fig. 2 depicts our StructTuning framework. For a limited set of curated domain corpora (typically textbooks), we first develop an automatic approach to extract the knowledge structure, and associate the training chunks to the extracted knowledge structure and points (Sec. 3.1). Then, we design a two-stage training strategy to inject the highly structured domain knowledge into language models by mimicking the human education process, comprising the SCPT (Sec. 3.2) and SSFT (Sec. 3.3) techniques. The details are introduced as follows.
@@ 3.1 Automatic Extraction of Knowledge Structure
For web-crawled corpus, the meta-info of knowledge structures (e.g., the table content for a textbook) is usually inaccessible or inapplicable due to the data-parsing quality limitation, and all we have are those sequentially arranged text segments (e.g., page-by-page-chunked content). As shown in Fig. 2 (a), we aim to extract (or, recover) the knowledge structure from the raw corpus for subsequent domain knowledge injection, which is achieved by the following steps.
First, we use the spaCy library [^1] https://github.com/explosion/spaCy to split the content from a textbook at the sentence-level, and merge the sentences to form training chunks within a maximum size (e.g., 2048 tokens [35]). After that, we prompt the advanced Llama3-70B [2] model to summarize the title for each chunk, where the textual content with the abstractive title contributes to a “knowledge point”.
Then, we explore an automatic method to aggregate the sequential knowledge points and extract the inherent structure hierarchy by leveraging advanced language models. Particularly, we take the title list of those text chunks as the input (due to the limitation of context length) to instruct Llama3-70B [2] to identify the inherent structure within the title/chunk list, through in-context learning on the provided examples. The prompt template is displayed in Fig. 3, and the full version is in Appendix D.
Through the top-down structurization on the title list, we manage to extract the hierarchical knowledge structure from the original training corpus. Fig. 3 presents an example of the extracted structure, where we use the tree-like mindmap structure [46] to present the knowledge taxonomy from a textbook. The whole process does not involve human annotation, which reduces the cost and makes our method scalable for handling larger amounts of domain training corpora.
Figure 3: Left: prompt template to extract hierarchical knowledge structures from the given content. Right: example of representing the extracted knowledge structure by using a mindmap template.
@@ 3.2 Structure-aware Continual Pre-Training
In conventional knowledge injection methods, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, leading to the fact that models can only absorb domain knowledge that is emergent in the data diversity [33, 30, 35]. In this section, we present another solution to inject knowledge from limited pieces of textbooks by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
@@
We first transform the knowledge structure into natural languages using the same mindmap template in Fig. 3, and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy. We collected 20 diversified templates from GPT-4 [1] to bridge mindmap structures and training chunks, one of which is displayed in Fig. 4. The prepended mindmap, as well as the template, does not produce auto-regressive loss. Losses are only calculated in the content part. Formally, we turn the original language modeling in vanilla CPT to conditioned modeling [22] in our SCPT stage:
@@
where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability [5] on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap paths.
@@ 3.3 Structure-aware Supervised Fine-Tuning
In traditional knowledge injection paradigms, supervised fine-tuning serves as a tool to align the (continually) pre-trained models to interactive chatbots, e.g., through massive question-answering exercises [9, 35]. However, most QA data augmentation or generation strategies aim at enlarging the quantity and enhancing the diversity of training samples [49, 31, 28], which neglects the nature of the highly structured domain knowledge. Therefore, our structure-aware supervised fine-tuning (SSFT) technique focuses on eliciting models’ structured knowledge learned during the SCPT stage, adapting LLMs to interactive and reliable domain experts.
Fig. 2 (c) illustrates the idea of synthesizing SSFT samples guided by domain knowledge structures, as extracted in Sec. 3.1. Specifically, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap. For paths with only one knowledge point, we use the corresponding text content to prompt Llama3-70B [2] to generate knowledge-intensive question-answering pairs. For paths with two or more branches, we prompt Llama3-70B with the knowledge path and textual contents to synthesize 2- or multi-hop QA samples, which require specific reasoning along the knowledge structure to derive from questions to answers. Fig. 5 presents several examples.
@@
For every synthesized QA sample (${\bm{z}}$), we will prepend the relevant mindmap hierarchy to the answer, and add a CoT prompt in the question to construct another type of QA data (${\bm{z}}^{\prime}$) for SFT alignment. This design explicitly elicits the learned knowledge in models’ responses, teaching them how to apply the structured knowledge to address real-world problems. We use the two types of QA samples for training, as recommended by Qiu et al. [35]. During testing, we can either use the vanilla question as input to efficiently gather models’ answers, or take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.

### v2

# Structure-aware Domain Knowledge Injection for Large Language Models

@@ Abstract
This paper introduces a pioneering methodology, termed StructTuning, to efficiently transform foundation Large Language Models (LLMs) into domain specialists. It significantly reduces the training corpus requirement to a mere 0.3%, while achieving an impressive 50% of traditional knowledge injection performance. Our method is inspired by the educational processes of human students, particularly how structured domain knowledge from textbooks is assimilated and subsequently applied to tackle real-world challenges through specific exercises. Based on this, we propose a novel two-stage strategy for knowledge injection and alignment: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT). In the SCPT phase, we automatically extract the domain knowledge taxonomy and reorganize the training corpora, enabling LLMs to effectively link textual segments to targeted knowledge points within the taxonomy. In the SSFT phase, we explicitly prompt models to elucidate the underlying knowledge structure in their outputs, leveraging the structured domain insight to address practical problems. Our ultimate method has undergone extensive evaluations across model architectures and scales, using closed-book question-answering tasks on LongBench and MMedBench datasets. Remarkably, our method demonstrates the potential of comparable improvement against the state-of-the-art MMedLM2 on MMedBench, while significantly reducing the training costs to 5%. This breakthrough paves the way for scaling up our StructTuning for stronger domain-specific LLMs with comprehensive data utilization. Code is available at https://github.com/alibaba/struxgpt.
@@ 1 Introduction
Large language models (LLMs) have recently seen extensive deployment across various applications (Vaswani et al., 2017; Achiam et al., 2023; Jiang et al., 2023; Bi et al., 2024). When adapting foundational models (e.g., Llama series (Touvron et al., 2023a; b; Dubey et al., 2024)) for specialized AI assistants in distinct domains (Qiu et al., 2024; Guo et al., 2024), developers usually employ two techniques to enhance LLMs’ proficiency: retrieval-augmented generation (RAG) (Lewis et al., 2020) and domain knowledge injection (Gururangan et al., 2020). While RAG effectively utilizes an external knowledge base to augment information, the retrieval process’s inherent noise poses challenges to generating reliable responses, especially in scenarios requiring logical reasoning where there is a semantic gap between the user’s query and the knowledge base.(Zhang et al., 2023; Chen et al., 2023). Thus, another avenue of research focuses on injecting new knowledge into models via training techniques (Gu et al., 2021; Hu et al., 2021; Mecklenburg et al., 2024).
Continual pre-training (Sun et al., 2020; Ibrahim et al., 2024) has been preferred for integrating new, domain-specific knowledge into existing LLMs (Cui et al., 2023; Wang et al., 2023b; Qiu et al., 2024). Nevertheless, it often entails resource-intensive auto-regressive training on billions of tokens from the internet to learn fragmented knowledge points, rather than absorbing structured knowledge from a few domain-specific textbooks (Jin et al., 2020). For example, MMedLM (Qiu et al., 2024) curates 25.5B tokens to derive a medical model, and DeepSeek-Coder (Guo et al., 2024) uses 2T tokens for coding adaptation. The common failure to learn effectively from limited textbook content has been attributed to insufficient data diversity (Zhu & Li, 2023a), which however violates the observation during the human education process in Fig. 1: students gain knowledge by sequentially studying from textbooks, reviewing knowledge points and structures, and applying this knowledge through proper exercises. In this process, all the new data to learn are textbooks (structured content) and exercising examples (question-answering pairs), and students just adopt their world knowledge to memorize, understand, and apply the knowledge to become domain experts (Krathwohl, 2002; Yu et al., 2023).
@@
Inspired by this, we propose to inject the domain knowledge from textbooks into LLMs, as educating a human student, through a novel two-stage training strategy: Structure-aware Continual Pre-Training (SCPT) and Structure-aware Supervised Fine-Tuning (SSFT).
@@
In the SCPT stage, we argue that high-quality textbook data can adequately infuse the domain knowledge (Gunasekar et al., 2023), where the organization of training corpora is crucial. In conventional paradigms, as illustrated in Fig. 1, text corpora are simply concatenated and divided into chunks of 2048 (Qiu et al., 2024) or 4096 (Guo et al., 2024), while the inherent structure of the texts (e.g., catalogs of textbooks) is disregarded. Instead, we propose an automatic approach to maintain each chunk’s knowledge structure. We view each chunk as a knowledge point, and employ advanced LLMs to efficiently extract domain knowledge taxonomy from the corpus, bypassing the need for manual annotation. Subsequently, LLMs are trained to predict the textual content (corresponding to the knowledge point) under the condition of the knowledge path within the domain structure, linking individual training chunks with the entire knowledge architecture. Finally, models are asked to memorize the knowledge structure to review the whole domain knowledge system.
@@
In the SSFT stage, the goal shifts from knowledge injection to enabling LLMs to recall and utilize their acquired knowledge to tackle real-world challenges. We explicitly elicit knowledge paths in LLMs’ responses, as a beacon for models to targeted information retrieval or logical reasoning for reliable responses. To this end, we derive a scalable strategy to generate question-answer pairs as practice exercises by powerful LLMs such as GPT4 (Achiam et al., 2023) or LLaMA3 (Dubey et al., 2024). In the scenarios with existing QA pairs like MMedBench (Qiu et al., 2024), we retrieve related knowledge structure and content, instructing LLaMA3 to provide explanations from questions to answers based on the knowledge paths. For datasets lacking specific QA samples like LongBench (Bai et al., 2023), we randomly select knowledge paths from the domain taxonomy and prompt LLaMA3 to craft question-answer-explanation triplets for training exercises.
@@
Our ultimate approach, termed StructTuning, outperforms conventional methods in domain knowledge injection by emulating human learning processes through SCPT and SSFT phases. We extensively evaluate StructTuning’s effectiveness across different model architectures and sizes. For domain-adapted language models, we first examine their capability to recall the injected knowledge through open-ended QA on the LongBench (Bai et al., 2023) dataset, then assess their application of this knowledge in addressing real-world issues via multiple-choice QA on MMedBench (Qiu et al., 2024). Both evaluations underscore the superiority of StructTuning. Remarkably, we achieve a 50% improvement in knowledge injection compared to the SOTA MMedLM2 in the medical domain, using merely 0.3% of the training data requirement. And StructTuning illustrates the potential of comparable performance with only 5% of training costs. These findings reveal the scalability of our method for enhancing domain-specific AI assistants with further comprehensive data utilization.
@@
  • We conducted extensive investigations on our StructTuning strategy on various data and model settings, and comprehensively illustrate our superiority in knowledge injection.
@@ 2 Related Work
Domain Adaptation for Large Language Models. While pre-trained LLMs possess promising capabilities, their performance is often hampered by the scope and recency of their training data, which particularly affects smaller models in downstream applications (Zhao et al., 2023; Wang et al., 2023a). Continual Pre-Training (CPT) addresses this by perpetually updating a pre-trained model with domain-specific content (Sun et al., 2020; Xu et al., 2023b). To keep pace with the latest information, models can be fine-tuned with supervised instruction-response pairs (SFT), thus staying current with the advancing knowledge landscape (Mecklenburg et al., 2024; Qiu et al., 2024). Existing literature confirms that combining CPT and SFT is effective for LLMs to remain precise and up-to-date in dynamic fields like law (Cui et al., 2023; Nguyen, 2023) and medicine (Wang et al., 2023b; Qiu et al., 2024). Our study builds upon this CPT-SFT framework, innovating with SCPT-SSFT strategies to efficiently and effectively infuse domain knowledge with the inherent structure hierarchy.
Structure-aware Knowledge Aggregation. A branch of researchers follows the conventional paradigm to extract entity-relation-entity triplets from texts to construct knowledge graphs (Pan et al., 2024), to enhance LLMs’s factual knowledge and logical reasoning by feature aggregation (Liu et al., 2020; Zhang et al., 2022), prompt engineering (Wen et al., 2023; Wang et al., 2023c), training data synthesis (Tang et al., 2024), etc. In these cases, each node corresponds to either a specific entity or an abstract concept, lacking the capability to present an informative and self-contained knowledge point. Some works have recently related a piece of descriptive text to a knowledge point, and constructed the knowledge structure for LLMs’ retrieval-augmented generation (Sarthi et al., 2024; Dong et al., 2024), where the top-to-down retrieval provides precise information-seeking paths along the knowledge structure. In this paper, we extend the structure-aware knowledge aggregation to LLMs’ training phase, injecting the whole domain knowledge structure into LLMs’ by linking the training samples into corresponding knowledge points and reasoning paths.
Data Augmentation and Synthesis. Traditional methods aim to artificially expand the training dataset size (Xu et al., 2023a; Mukherjee et al., 2023) or generate entirely new samples that could help models learn better or adapt to specific tasks (Tang et al., 2024). Yet, they often overlook the structured nature of domain knowledge, and the aimlessly generated samples may also lack diversity (Ovadia et al., 2023; Mecklenburg et al., 2024), leading to potentially suboptimal training outcomes when applied for domain adaptations (Mecklenburg et al., 2024; Tang et al., 2024). By contrast, our SSFT design is an innovative departure to address the challenge of retaining and utilizing the structured knowledge inherent in domain-specific content.
@@ 3 Methodology
Fig. 2 depicts our StructTuning methodology to inject domain knowledge into pre-trained LLMs using the inherent knowledge structure. With curated domain corpora (typically a few textbooks), we first develop an automatic approach to extract the knowledge structure, and associate text chunks to corresponding knowledge paths and points (Sec. 3.1). Then, we design a two-stage training strategy to inject the highly structured domain knowledge into language models by mimicking the human education process, comprising the SCPT (Sec. 3.2) and SSFT (Sec. 3.3) techniques.
@@ 3.1 Automatic Extraction of Knowledge Structure
For web-crawled corpus, previous data pre-processing focuses on quality assessment for individual documents (Bi et al., 2024), while the meta-info of knowledge structures (e.g., the table content for a textbook) is usually neglected or filtered out, and all we have are those sequentially arranged text segments (e.g., page-by-page-chunked content). As shown in Fig. 2 (a), we aim to extract (or, recover) the knowledge structure from the raw corpus for subsequent domain knowledge injection.
First, we use spaCy [^1] https://github.com/explosion/spaCy to split the content from a textbook at the paragraph-level, and merge the sentences to form training chunks within a maximum size (e.g., 2048 tokens (Qiu et al., 2024)). After that, we prompt the advanced Llama3-70B (Dubey et al., 2024) model to summarize the title for each chunk, where the textual content with the abstractive title jointly contributes to a “knowledge point”.
Then, we automatically aggregate knowledge points and extract the inherent structure hierarchy by leveraging advanced language models. Inspired by Liu et al. (2024a), we take the title list to instruct a specifically developed 7B model to identify the inherent knowledge structure within the text chunks. A prompt example is displayed in Fig. 3, and the detailed implementation is presented in Sec. B.1. In particular, Sec. B.1 and Sec. B.2 verify that our specialized 7B model can identify sufficiently precise knowledge structure for effective and efficient domain adaptation, as more powerful LLMs like LLaMA3-70B (Dubey et al., 2024) and GPT-3.5 (Brown et al., 2020) cannot bring significant enhancement while largely increase the inference costs.
Fig. 3 presents an example of the extracted structure, where we use the tree-like mindmap structure (Wen et al., 2023) to present the knowledge taxonomy from a textbook. And Fig. A1 showcases how to deal with non-textbook data. The whole process does not involve human annotation, which reduces the cost and makes our method scalable for larger domain training corpora.
Figure 3: Left: prompt template to extract hierarchical knowledge structures from the given content. Right: example of the extracted knowledge structure presented by a mindmap format.
@@ 3.2 Structure-aware Continual Pre-Training
In conventional knowledge injection methods, training corpora are randomly concatenated and chunked into text segments without distinguishing the original content, leading to the fact that models can only absorb domain knowledge that is emergent in the data diversity (Ovadia et al., 2023; Mecklenburg et al., 2024; Qiu et al., 2024). In this section, we present another solution to inject knowledge from limited pieces of textbooks by leveraging the highly abstractive and exhaustive domain knowledge structures for continual pre-training.
@@
We first transform the knowledge structure into natural languages using the same mindmap template in Fig. 3, and prepend it to each training chunk, forcing LLMs to memorize the textual content (knowledge points) in the condition of the associated knowledge path in the structure hierarchy. We collected 20 diversified templates from GPT-4 (Achiam et al., 2023) to bridge mindmap structures and training chunks, one of which is displayed in Fig. 4. The prepended mindmap, as well as the template, does not produce auto-regressive loss. Losses are only calculated in the content part. Formally, we turn the original language modeling in vanilla CPT to conditioned modeling (Keskar et al., 2019) in our SCPT stage:
@@
where $p({\bm{x}}^{k})$ models the probability distribution for the $k$-th chunk ${\bm{x}}^{k}=(x^{k}_{1},\cdots,x^{k}_{n})$ via the chain rule of probability (Bengio et al., 2000) on each token $x^{k}_{i}$, and ${\bm{s}}^{k}$ denotes the associated knowledge mindmap paths.
@@ 3.3 Structure-aware Supervised Fine-Tuning
In traditional knowledge injection paradigms, supervised fine-tuning serves as a tool to align the (continually) pre-trained models to interactive chatbots, e.g., through massive question-answering exercises (Cui et al., 2023; Qiu et al., 2024). However, most QA data augmentation or generation strategies aim at enlarging the quantity and enhancing the diversity of training samples (Xu et al., 2023a; Mukherjee et al., 2023; Liu et al., 2024b), which neglects the nature of the highly structured domain knowledge. Therefore, our structure-aware supervised fine-tuning (SSFT) technique focuses on eliciting models’ structured knowledge learned during the SCPT stage, adapting LLMs to interactive and reliable domain experts.
Fig. 2 (c) illustrates the idea of synthesizing SSFT samples guided by domain knowledge structures. Specifically, we use the random walk algorithm to create knowledge paths with $1$ to $l$ branches in the original mindmap. For paths linking to a single knowledge point, we use the corresponding text content to prompt Llama3-70B (Dubey et al., 2024) to generate knowledge-intensive question-answering pairs. For paths with two or more branches, we prompt Llama3-70B with the knowledge path and textual contents to synthesize 2- or multi-hop QA samples, which require specific reasoning along the knowledge structure to derive from questions to answers. Fig. 5 presents several examples.
@@
For every synthesized QA sample (${\bm{z}}$), we will prepend the relevant mindmap hierarchy to the answer, and add a CoT prompt in the question to construct another type of QA data (${\bm{z}}^{\prime}$) for SFT alignment. This design explicitly elicits the learned knowledge in models’ responses, teaching them how to apply the structured knowledge to address real-world problems. We use the two types of QA samples for training, as recommended by Qiu et al. (2024). During testing, we can either use the vanilla question as input to efficiently gather models’ answers, or take the CoT prompt to probe to what extent LLMs can memorize and leverage the injected knowledge to answer the questions.